# On schurity of the dihedral group $\boldsymbol{D_{2p}}$

### v3

@@ 4. $S$-rings over $D_{2p}$
It should be mentioned that every $S$-ring over $C_{2p}$ is of rank $2$, or cyclotomic, or a wreath product of two $S$-rings by [9, Theorem 4.1-4.2] (see also [20, 21]). The proof of Theorem 4.1 will be given in the end of the section. We say that a subset $X$ of $G$ is mixed if $X\cap A\neq\varnothing$ and $X\cap Ab\neq\varnothing$; otherwise we say that $X$ is nonmixed. Throughout this section, $\mathcal{A}$ is an $S$-ring over $G$. The next lemma immediately follows from the fact that $G$ is a strong B-group (see [34, Theorem 26.6(i)]).
@@
If $G$ has an $\mathcal{A}$-subgroup $L$ of order $2$, then $\mathcal{A}$ is cyclotomic or isomorphic to an $S$-ring over $C_{2p}$.
@@
Let $L=\{e,s\}$. Then $\{s\}\in\mathcal{S}(\mathcal{A})$. Suppose that $A$ is an $\mathcal{A}$-subgroup. Lemma 2.1 implies that $\mathcal{A}_{A}=\operatorname{Cyc}(K_{0},A)$ for some $K_{0}\leq\operatorname{Aut}(A)$. Every basic set of $\mathcal{A}$ outside $A\cup L$ is of the form $Xs$, where $X\in\mathcal{S}(\mathcal{A}_{A})=\operatorname{Orb}(K_{0},A)$. Therefore $\mathcal{A}=\operatorname{Cyc}(K,G)$, where $K\leq\operatorname{Aut}(G)$ is such that the restriction of $K$ on $A$ is equal to $K_{0}$ and $K$ fixes $s$. Further we assume that $A$ is not an $\mathcal{A}$-subgroup.
@@
Due to the above paragraph, we may assume that every basic set $Y$ of $\mathcal{A}$ outside $L$ is mixed and hence of the form $Y=Y_{0}\cup Y_{1}s$, where $Y_{0}$ and $Y_{1}$ are nonempty subsets of $A$. Note that $Y_{1}s=(Y_{1}s)^{-1}$. So $Y=Y^{-1}$ and $Y_{0}=Y_{0}^{-1}$. Since $\{s\}\in\mathcal{S}(\mathcal{A})$, the set $Ys=Y_{0}s\cup Y_{1}$ is basic. Moreover, $Ys=(Ys)^{-1}$ and hence $Y_{1}=Y_{1}^{-1}$. Therefore $\sigma\in\operatorname{Aut}(G)$ such that $\sigma:(a,s)\mapsto(a^{-1},s)$ is a Cayley automorphism of $\mathcal{A}$. One can verify straightforwardly that $|\sigma b_{r}|=2$ and $a_{r}\sigma b_{r}=\sigma b_{r}a_{r}$. This implies that the group $K=\langle a_{r},\sigma b_{r}\rangle$ is isomorphic to $C_{2p}$. It is easy to see that $K$ is transitive on $G$ and hence regular. Thus, $\operatorname{Aut}(\mathcal{A})$ has a regular subgroup isomorphic to $C_{2p}$. This yields that $\mathcal{A}$ is isomorphic to an $S$-ring over $C_{2p}$. ∎
@@
[…]
In this case $Ab$ is an $\mathcal{A}$-set. Due to Lemma 2.1, we have $\mathcal{A}_{A}=\operatorname{Cyc}(K,A)$ for some $K\leq\operatorname{Aut}(A)\cong C_{p-1}$. Let $|K|=m$. Clearly, $m$ divides $p-1$ and every nontrivial basic set of $\mathcal{A}_{A}$ is of size $m$.
@@
Obviously, $|Y|\leq|Ab|=p$. If $|Y|=1$ ($|Y|=p-1$, respectively), then $\langle Y\rangle$ ($\langle Ab\setminus Y\rangle$, respectively) is an $\mathcal{A}$-subgroup of order $2$, a contradiction to the assumption that $A$ is the unique $\mathcal{A}$-subgroup. ∎
@@
Let $r$ be a prime and $s$ a positive integer. Suppose that $r^{s}$ divides $m$. Then there exists a unique $Y\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $r^{s}$ does not divide $|Y|$. Moreover, $|Y|\equiv 1~{}\mod~{}r^{s}$.
@@
Let $Y_{1},Y_{2}\in\mathcal{S}(\mathcal{A})_{Ab}$. Then $Y_{1}=Y_{1}^{-1}$ and $Y_{2}=Y_{2}^{-1}$. If $Y_{1}=Y_{2}$, then
$\underline{Y_{1}}\cdot\underline{Y_{1}}=|Y_{1}|e+\sum_{\begin{subarray}{c}X\in\mathcal{S}(\mathcal{A})_{A}\\ X\neq\{e\}\end{subarray}}c_{Y_{1}Y_{2}}^{X}\underline{X},$
and if $Y_{1}\neq Y_{2}$, then
@@
Observe that $\sum\limits_{Y\in\mathcal{S}(\mathcal{A})_{Ab}}|Y|=|Ab|=p$. This implies that there exists $Y\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $r$ does not divide $|Y|$; for otherwise $r$ divides $p$ which is impossible. The second part of Eq. (6) yields that $|Y||Z|$ is divided by $r^{s}$ and hence $|Z|$ is divided by $r^{s}$ for every $Z\in\mathcal{S}(\mathcal{A})_{Ab}$ with $Z\neq Y$. From the first part of Eq. (6) it follows that $|Y|-1$ is divided by $r^{s}$ or, equivalently, $|Y|\equiv 1~{}\mod~{}r^{s}$. ∎
@@
From Lemma 4.5 it follows that $|Z|$ is divided by $m$ for every $Z\in\mathcal{S}(\mathcal{A})_{Ab}$ with $Z\neq Y$. Note that $|Y|\geq 2$ by Lemma 4.4. Let $y_{1},y_{2}\in Yb$ such that $y_{1}\neq y_{2}$ and $X$ a nontrivial basic set of $\mathcal{A}_{A}$ such that $y_{2}y_{1}^{-1}\in X$. Since $X\subseteq A$ and $Y\subseteq Ab$, we have
@@
Counting the number of elements in the left-hand side and the right-hand side of the above equality and taking into account that $|X|=m$ and $|Z|$ is divided by $m$ for every $Z\in\mathcal{S}(\mathcal{A})_{Ab}$ distinct from $Y$, we conclude that
@@
for some nonnegative integer $\alpha$. Since $|Y|$ and $m$ are coprime, the number $c_{XY}^{Y}$ is divided by $m$ due to Eq. (8). Observe that $c_{XY}^{Y}\geq 1$ because $y_{2}y_{1}^{-1}\in X$ and the element $y_{2}b$ enters the element $\underline{X}\cdot\underline{Y}$ with nonzero coefficient. So $c_{XY}^{Y}\geq m$. Now Eq. (8) yields that $c_{XY}^{Y}=m$ and $\alpha=0$. Therefore from Eq. (7) it follows that $\underline{X}\cdot\underline{Y}=|X|\underline{Y}$ and hence $X\leq\operatorname{rad}(Y)$. Thus, $A=\langle X\rangle\leq\operatorname{rad}(Y)$ and $Y=Ab$. The latter means that $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$. ∎
@@
If $m=1$, then we are done by Lemma 4.6. If $m>1$, then there exists the unique $Y\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $\operatorname{GCD}(|Y|,m)=1$ by Lemma 4.5 because $m$ is a prime power. Again, we are done by Lemma 4.6. ∎
@@
Note that $m$ divides $p-1$. If $m\in\{1,2,4,q\}$, then we are done by Lemma 4.7. Since $m\neq p-1$, it remains to consider the case $r=4$ and $m=2q$. In this case, $\operatorname{rk}(\mathcal{A}_{A})=3$. Let $X_{1}$ and $X_{2}$ be nontrivial basic sets of $\mathcal{A}_{A}$. Clearly, $|X_{1}|=|X_{2}|=2q$. Since $m$ is even, we have $X_{1}=X_{1}^{-1}$ and $X_{2}=X_{2}^{-1}$. Thus, $\mathcal{A}$ is symmetric and hence commutative.
Lemma 4.5 implies that there exist a unique $Y_{1}\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $|Y_{1}|$ is not divided by $q$ and a unique $Y_{2}\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $|Y_{2}|$ is not divided by $2$. If $Y_{1}=Y_{2}$, then $\operatorname{GCD}(|Y_{1}|,m)=1$ and we are done by Lemma 4.6.
Suppose that $Y_{1}\neq Y_{2}$. From Lemma 4.5 it follows that $|Y_{1}|\equiv 1\mod q$. Since $Y_{2}$ is a unique basic set from $\mathcal{S}(\mathcal{A})_{Ab}$ of odd cardinality and $Y_{1}\neq Y_{2}$, we conclude that $|Y_{1}|$ is divided by $2$. Therefore
@@
Let $|Y_{1}|=3q+1$. Since $|Y_{2}|$ is divided by $q$, we have $\mathcal{S}(\mathcal{A})_{Ab}=\{Y_{1},Y_{2}\}$ and $|Y_{2}|=q$ in this case. Note that
@@
Since $|Y_{1}|$ is not divided by $q$ and $|X_{1}|=|X_{2}|=2q$, the above equalities yield that $c_{Y_{1}X_{1}}^{Y_{1}}$ and $c_{Y_{1}X_{2}}^{Y_{1}}$ are divided by $q$. If $|Y_{1}|=3q+1$, then $c_{Y_{1}X_{1}}^{Y_{1}}\geq 2q$ or $c_{Y_{1}X_{2}}^{Y_{1}}\geq 2q$ by Eq. (10). The latter means that $Y_{1}X_{1}=Y_{1}$ or $Y_{1}X_{2}=Y_{1}$. In both cases $\operatorname{rad}(Y_{1})$ is nontrivial and hence $Y_{1}=Ab$, a contradiction to $Y_{1}\neq Y_{2}$.
@@
for any $z,z^{\prime}\in Z_{1}$. Clearly, $|Z_{1}|=|Y_{1}|=q+1\geq 8$. Let $z,z^{\prime},z^{\prime\prime}\in Z_{1}$ be pairwise distinct. Since $p\geq 29$,
@@
by [4, Corollary 0.2]. Using Eqs. (11), (12), (13), we obtain
@@
$-|X_{1}z\cap X_{1}z^{\prime}|-|X_{1}z\cap X_{1}z^{\prime}|-|X_{1}z\cap X_{1}z^{\prime}|+|Z_{1}\cap X_{1}z\cap X_{1}z^{\prime}|+|Z_{1}\cap X_{1}z\cap X_{1}z^{\prime\prime}|+$
@@
Now suppose that there is no an $\mathcal{A}$-subgroup of order $2$. Since $\mathcal{A}$ is imprimitive, $A$ is a unique $\mathcal{A}$-subgroup. If $p$ is a Fermat prime, then $m$ is a $2$-power and $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$ by Lemma 4.7 and Statement $(4)$ of Theorem 4.1 holds. Let $p=rq+1$, where $q$ is a prime and $r\in\{2,4\}$. We may assume that $q$ is odd because otherwise $p$ is a Fermat prime and we are done. If $m\neq p-1$, then $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$ by Lemma 4.8 and Statement $(4)$ of Theorem 4.1 holds.
@@
Let $p$ be a Fermat prime or $p=4q+1$, where $q$ is prime. Let us prove that $\mathcal{A}$ is schurian. One of the statements of Theorem 4.1 holds for $\mathcal{A}$. If Statement $(1)$ or $(2)$ holds, then $\mathcal{A}$ is obviously schurian. If Statement $(3)$ holds, then $\mathcal{A}$ is schurian because $C_{2p}$ is a Schur group by [10, Theorem 1.1]. If Statement $(4)$ holds, then $\mathcal{A}$ is schurian by Lemma 2.1 and Lemma 2.2. If Statement $(5)$ holds, then $p=13$ and $\mathcal{A}$ is schurian by computer calculations [36]. ∎

### v4

@@ 7. Schurity of $D_{2p}$
It should be mentioned that every $S$-ring over $C_{2p}$ is of rank $2$, or cyclotomic, or a wreath product of two $S$-rings by [13, Theorem 4.1-4.2] (see also [25, 26]). The proof of Theorem 7.1 will be given in the end of the section. We say that a subset $X$ of $G$ is mixed if $X\cap A\neq\varnothing$ and $X\cap Ab\neq\varnothing$; otherwise we say that $X$ is nonmixed. Throughout this section, $\mathcal{A}$ is an $S$-ring over $G$. The next lemma immediately follows from the fact that $G$ is a strong B-group (see [42, Theorem 26.6(i)]).
@@
If $G$ has an $\mathcal{A}$-subgroup of order $2$, then $\mathcal{A}$ is cyclotomic or isomorphic to an $S$-ring over $C_{2p}$.
@@
Let $L=\{e,s\}$ be an $\mathcal{A}$-subgroup. Then $\{s\}\in\mathcal{S}(\mathcal{A})$. Suppose that $A$ is an $\mathcal{A}$-subgroup. Lemma 2.3(1) implies that $\mathcal{A}_{A}=\operatorname{Cyc}(K_{0},A)$ for some $K_{0}\leq\operatorname{Aut}(A)$. Every basic set of $\mathcal{A}$ outside $A\cup L$ is of the form $Xs$, where $X\in\mathcal{S}(\mathcal{A}_{A})=\operatorname{Orb}(K_{0},A)$. Therefore $\mathcal{A}=\operatorname{Cyc}(K,G)$, where $K\leq\operatorname{Aut}(G)$ is such that the restriction of $K$ on $A$ is equal to $K_{0}$ and $K$ fixes $s$. Further we assume that $A$ is not an $\mathcal{A}$-subgroup.
@@
Due to the above paragraph, we may assume that every basic set $Y$ of $\mathcal{A}$ outside $L$ is mixed and hence of the form $Y=Y_{0}\cup Y_{1}s$, where $Y_{0}$ and $Y_{1}$ are nonempty subsets of $A$. Note that $Y_{1}s=(Y_{1}s)^{-1}$. So $Y=Y^{-1}$ and $Y_{0}=Y_{0}^{-1}$. Since $\{s\}\in\mathcal{S}(\mathcal{A})$, the set $Ys=Y_{0}s\cup Y_{1}$ is basic. Moreover, $Ys=(Ys)^{-1}$ and hence $Y_{1}=Y_{1}^{-1}$. Therefore $\sigma\in\operatorname{Aut}(G)$ such that $a^{\sigma}=a^{-1}$ and $s^{\sigma}=s$ is a Cayley automorphism of $\mathcal{A}$. One can verify straightforwardly that $|\sigma s_{r}|=2$ and $a_{r}\sigma s_{r}=\sigma s_{r}a_{r}$. This implies that the group $R=\langle a_{r},\sigma s_{r}\rangle$ is isomorphic to $C_{2p}$. It is easy to see that $R$ is transitive on $G$ and hence regular. Thus, $\operatorname{Aut}(\mathcal{A})$ has a regular subgroup isomorphic to $C_{2p}$. This yields that $\mathcal{A}$ is isomorphic to an $S$-ring over $C_{2p}$. ∎
@@
$A~{}\text{is a unique}~{}\mathcal{A}\text{-subgroup of}~{}G$
In this case, $Ab$ is an $\mathcal{A}$-set. Due to Lemma 2.3(1), we have $\mathcal{A}_{A}=\operatorname{Cyc}(K,A)$ for some $K\leq\operatorname{Aut}(A)\cong C_{p-1}$. Let $|K|=m$. Clearly, $m$ divides $p-1$ and every nontrivial basic set of $\mathcal{A}_{A}$ is of size $m$.
@@
Obviously, $|Y|\leq|Ab|=p$. If $|Y|=1$ ($|Y|=p-1$, respectively), then $\langle Y\rangle$ ($\langle Ab\setminus Y\rangle$, respectively) is an $\mathcal{A}$-subgroup of order $2$, a contradiction to the assumption that $A$ is a unique $\mathcal{A}$-subgroup. ∎
@@
Let $r$ be a prime and $s$ a positive integer. Suppose that $r^{s}$ divides $m$. Then there exists a unique $Y\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $r^{s}$ does not divide $|Y|$. Moreover, $|Y|\equiv 1~{}\bmod r^{s}$.
@@
Let $Y_{1},Y_{2}\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $Y_{1}\neq Y_{2}$. Clearly, $Y_{1}=Y_{1}^{-1}$ and $Y_{2}=Y_{2}^{-1}$. One can see that
$\underline{Y_{1}}\cdot\underline{Y_{1}}=|Y_{1}|e+\sum_{\begin{subarray}{c}X\in\mathcal{S}(\mathcal{A})_{A}\\ X\neq\{e\}\end{subarray}}c_{Y_{1}Y_{1}}^{X}\underline{X},$
and
@@
Observe that $\sum\limits_{Y\in\mathcal{S}(\mathcal{A})_{Ab}}|Y|=|Ab|=p$. This implies that there exists $Y\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $r$ does not divide $|Y|$; for otherwise $r$ divides $p$ which is impossible. The second part of Eq. (15) yields that $|Y||Z|$ is divisible by $r^{s}$ and hence $|Z|$ is divisible by $r^{s}$ for every $Z\in\mathcal{S}(\mathcal{A})_{Ab}$ with $Z\neq Y$. From the first part of Eq. (15) it follows that $|Y|-1$ is divisible by $r^{s}$ or, equivalently, $|Y|\equiv 1~{}\bmod r^{s}$. ∎
@@
From Lemma 7.5 it follows that $|Z|$ is divisible by $m$ for every $Z\in\mathcal{S}(\mathcal{A})_{Ab}$ with $Z\neq Y$. Note that $|Y|\geq 2$ by Lemma 7.4. Let $y_{1},y_{2}\in Yb$ such that $y_{1}\neq y_{2}$ and $X$ a nontrivial basic set of $\mathcal{A}_{A}$ such that $y_{2}y_{1}^{-1}\in X$. Since $X\subseteq A$ and $Y\subseteq Ab$, we have
@@
Counting the number of elements in the left-hand side and the right-hand side of the above equality and taking into account that $|X|=m$ and $|Z|$ is divisible by $m$ for every $Z\in\mathcal{S}(\mathcal{A})_{Ab}$ distinct from $Y$, we conclude that
@@
for some nonnegative integer $\alpha$. Since $|Y|$ and $m$ are coprime, the number $c_{XY}^{Y}$ is divisible by $m$ due to Eq. (17). Observe that $c_{XY}^{Y}\geq 1$ because $y_{2}y_{1}^{-1}\in X$ and the element $y_{2}b$ enters the element $\underline{X}\cdot\underline{Y}$ with nonzero coefficient. So $c_{XY}^{Y}\geq m$. Now Eq. (17) yields that $c_{XY}^{Y}=m$ and $\alpha=0$. Therefore from Eq. (16) it follows that $\underline{X}\cdot\underline{Y}=|X|\underline{Y}$ and hence $X\leq\operatorname{rad}(Y)$. Thus, $A=\langle X\rangle\leq\operatorname{rad}(Y)$ and $Y=Ab$. The latter means that $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$. ∎
@@
If $m=1$, then we are done by Lemma 7.6. If $m>1$, then there exists a unique $Y\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $\operatorname{GCD}(|Y|,m)=1$ by Lemma 7.5. Again, we are done by Lemma 7.6. ∎
@@
The condition of the lemma implies that $q$ is odd. Note that $m$ divides $p-1$. If $m\in\{1,2,4,q\}$, then we are done by Lemma 7.7. Since $m\neq p-1$, it remains to consider the case $r=4$ and $m=2q$. In this case, $\operatorname{rk}(\mathcal{A}_{A})=3$. Let $X_{1}$ and $X_{2}$ be nontrivial basic sets of $\mathcal{A}_{A}$. Clearly, $|X_{1}|=|X_{2}|=2q$. Since $m$ is even, $K$ contains an automorphism of order $2$ and consequently $X_{1}=X_{1}^{-1}$ and $X_{2}=X_{2}^{-1}$. Thus, $\mathcal{A}$ is symmetric and hence commutative.
Lemma 7.5 implies that there exist a unique $Y_{1}\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $|Y_{1}|$ is not divisible by $q$ and a unique $Y_{2}\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $|Y_{2}|$ is not divisible by $2$. If $Y_{1}=Y_{2}$, then $\operatorname{GCD}(|Y_{1}|,m)=1$ and we are done by Lemma 7.6.
Suppose that $Y_{1}\neq Y_{2}$. From Lemma 7.5 it follows that $|Y_{1}|\equiv 1\bmod q$. Since $Y_{2}$ is a unique basic set from $\mathcal{S}(\mathcal{A})_{Ab}$ of odd cardinality and $Y_{1}\neq Y_{2}$, we conclude that $|Y_{1}|$ is divisible by $2$. Therefore
@@
Note that
@@
Since $|Y_{1}|$ is not divisible by $q$ and $|X_{1}|=|X_{2}|=2q$, the above equalities yield that $c_{Y_{1}X_{1}}^{Y_{1}}$ and $c_{Y_{1}X_{2}}^{Y_{1}}$ are divisible by $q$. If $|Y_{1}|=3q+1$, then $c_{Y_{1}X_{1}}^{Y_{1}}\geq 2q$ or $c_{Y_{1}X_{2}}^{Y_{1}}\geq 2q$ by Eq. (19). The latter means that $Y_{1}X_{1}=Y_{1}$ or $Y_{1}X_{2}=Y_{1}$. In both cases, $\operatorname{rad}(Y_{1})$ is nontrivial and hence $Y_{1}=Ab$, a contradiction to $Y_{1}\neq Y_{2}$.
@@
for all $z,z^{\prime}\in Z_{1}$. Clearly, $|Z_{1}|=|Y_{1}|=q+1\geq 8$. Let $z,z^{\prime},z^{\prime\prime}\in Z_{1}$ be pairwise distinct. Since $p\geq 29$,
@@
by [5, Corollary 0.2] stating that in this case any three distinct vertices in a Paley graph have a common neighbor. Using Eqs. (20), (21), (22), we obtain
@@
$-|X_{1}z\cap X_{1}z^{\prime}|-|X_{1}z\cap X_{1}z^{\prime}|-|X_{1}z^{\prime}\cap X_{1}z^{\prime\prime}|+|Z_{1}\cap X_{1}z\cap X_{1}z^{\prime}|+|Z_{1}\cap X_{1}z\cap X_{1}z^{\prime\prime}|+$
@@
Lemma 7.9.
If $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$, then $\mathcal{A}$ is isomorphic to an $S$-ring over $C_{2p}$.
Proof.
Let $\sigma\in\operatorname{Aut}(G)$ such that $a^{\sigma}=a$ and $b^{\sigma}=ab$. It is easy to check that the group $R=\langle b_{r}\sigma\rangle$ is a regular subgroup of $\operatorname{Aut}(\mathcal{A})$ isomorphic to $C_{2p}$. ∎
@@
Now suppose that there is no an $\mathcal{A}$-subgroup of order $2$. Since $\mathcal{A}$ is imprimitive, $A$ is a unique $\mathcal{A}$-subgroup. In view of Lemma 7.9, to prove that Statement $(3)$ of Theorem 7.1 holds for $\mathcal{A}$, it suffices to verify that $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$. If $p$ is a Fermat prime, then $m$ is a $2$-power and $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$ by Lemma 7.7. Let $p=rq+1$, where $q$ is a prime and $r\in\{2,4\}$. We may assume that $q$ is odd because otherwise $p$ is a Fermat prime and we are done. If $m\neq p-1$, then $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$ by Lemma 7.8.
@@
Let $p$ be a Fermat prime or $p=4q+1$, where $q$ is prime. Let us prove that $\mathcal{A}$ is schurian. One of the statements of Theorem 7.1 holds for $\mathcal{A}$. If Statement $(1)$ or $(2)$ holds, then $\mathcal{A}$ is obviously schurian. If Statement $(3)$ holds, then $\mathcal{A}$ is schurian because $C_{2p}$ is a Schur group by [14, Theorem 1.1]. If Statement $(4)$ holds, then $p\in\{5,13\}$ and $\mathcal{A}$ is schurian by computer calculations [44]. ∎